OpenStudy (pphalke):

Michelle is 25 years old and wants to retire at the age of 50. if she invests $60,000 inheritance at 7% compounded semi annually how much will she have when she turns 50

OpenStudy (pphalke):

@Mehek14 @SnuggieLad

satellite73 (satellite73):

\[60,000\times (1+\frac{.07}{2})^{50}\] and a calculator

OpenStudy (pphalke):

i don't have a right calculator for this

satellite73 (satellite73):

your kidding right?

satellite73 (satellite73):

you are on a computer?

satellite73 (satellite73):

http://www.wolframalpha.com/input/?i=60,000 \times+%281%2B\frac{.07}{2}%29^{50}

OpenStudy (pphalke):

yea my hand held isn't a scientific one

satellite73 (satellite73):

copy and paste

OpenStudy (sshayer):

in the chrome ,write scientific computer and click,you will get.

Vocaloid (vocaloid):

pro tip: for simple calculations like this you can just type it into google

Vocaloid (vocaloid):

type in 60,000 * (1+0.07/2)^50 into the google search bar